# Gated Rotary-Enhanced Linear Attention for Long-term Sequential Recommendation

### v1

@@ 3.4. Prediction Layer and Training
(24) $\hat{\mathcal{Y}}=\operatorname{Softmax}(\boldsymbol{E}^{\top}\boldsymbol{H}_{L}).$
Hence, following previous works (Kang and McAuley, 2018; Liu et al., 2024b), we expect to minimize the cross-entropy between the predicted recommendation results $\hat{\mathcal{Y}}$, and the ground truth items $\mathcal{Y}$:
(25) $\mathcal{L}_{CE}\left(\mathcal{Y},\hat{\mathcal{Y}}\right)=\mathcal{Y}\log\left(\hat{\mathcal{Y}}\right)+\left(1-\mathcal{Y}\right)\left(1-\log\left(\hat{\mathcal{Y}}\right)\right).$
3.5. Complexity Analysis
@@
4. Experiments
Table 3. Statistics of the experimental datasets.
@@ 4. Experiments
4.1. Experimental Settings
4.1.1. Datasets
We conduct extensive experiments and verify the effectiveness of our approach on four real-world datasets, including MovieLens-1M, MovieLens-20M, MovieLens-32M  [^2] https://grouplens.org/datasets/movielens/ (dubbed as ML-1M, ML-20M, ML-32M) and Netflix  [^3] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data. The dataset statistics are shown in Table 3, where Avg. UA and Avg. IA denote the average actions of users and items. Notably, for MovieLens datasets, we filter out users and items with fewer than 5 interactions. Netflix is a huge-scale dataset, containing approximately 100 million interactions and five thousand Avg. IA. To simplify the training process, we filter out users with fewer than 50 interactions while retaining items with interactions ranging from 50 to 15,000 for this dataset.
4.1.2. Baseline Models
To verify the effectiveness and efficiency of RecGRELA, we make comparisons between RecGRELA and strong RNN-, Attention-, and Mamba-based baselines:
  • LinRec (Liu et al., 2023): This model reduces the computational costs via linear attention in transformer-based models. We choose SASRec (Kang and McAuley, 2018) as the backbone of LinRec.
  • LRURec (Yue et al., 2024): This method is based on linear recurrent units (Orvieto et al., 2023) for sequential recommendation.
  • Mamba4Rec (Liu et al., 2024b): This is the first Mamba-based method, which explores applying the Mamba architecture in sequential recommendation tasks.
  • EchoMamba4Rec (Wang et al., 2024): This method processes sequences both forward and backward to enhance the sequential representation with Mamba and frequency domain filter.
  • EulerFormer (Tian et al., 2024): This method proposes a transformer variant with complex vectors based on Euler’s formula and a contrastive learning method to improve contextual representations.
  • RecBLR (Liu et al., 2024a): This model shows a behavior-dependent linear recurrent unit to enhance user behavior modeling with a hardware-aware parallel scan approach.
  • SIGMA (Liu et al., 2025): This model employs a flipped Mamba to build a bidirectional framework to improve short sequence representations and contextual modeling.
Table 4. Performance comparisons on ML-1M, ML-20M, ML-32M, and Netflix datasets. Best and second-best results per row are shown in bold and underlined format, respectively. ”Improv.” denotes the percentage improvement of the best-performing model over the second-best for each metric.
@@ 4.1.2. Baseline Models
Figure 3. RecGRELA vs. baselines in terms of the GPU memory and FLOPs of the training stage on the ML-1M dataset.
4.1.3. Evaluation Protocols
@@
4.1.4. Implementation Details
We utilize the RecBole library  [^4] https://github.com/RUCAIBox/RecBole (Zhao et al., 2021; Xu et al., 2023), a unified framework for implementing sequential recommendations. To ensure a fair comparison, we fix the embedding size to 64, the hidden size of the MLP to 256, and the training batch size to 2048 for all the models. We directly adopt the values reported in their original manuscripts for the remaining hyperparameters of the baseline models. The number of blocks $L$ is set to 4. The number of attention heads $h$ is set to 4. The kernel size for 1D causal convolution is 4. We set the Dropout and Droppath rate to 0.1. We use Adam (Kingma, 2014) with a learning rate of 0.001 for optimizing our model. The maximum sequence length $N$ is set to 200 for all the datasets. The early stopping epochs are 20 for ML-1M and 10 for other datasets. All the experiments are conducted on a single NVIDIA RTX4090.
4.2. Overall Performance
Table 4 reports the recommendation performance of RecGRELA compared to seven baselines on four datasets. From the table, we find several observations:
@@ 4.3. Efficiency Analysis
Table 5. Sensitivity analysis on different numbers of layers. The sequence length is set to 200. Time indicates the training time per epoch, and Memory indicates the GPU memory.
@@
Figure 4. The effectiveness of each variant of RecGRELA on ML-1M.
4.4. Ablation Study
4.4.1. Core Components
@@
Figure 5. The performance of different position encodings and activation functions.
4.4.2. Position Encoding and Activation Function
@@
Figure 6. Impact of different stacked layers on computational costs. We set the sequence length to 100 to avoid GPU memory errors.
4.5. Sensitivity Analysis
@@ 4.5.1. The Number of Layers
Figure 7. Impact of sequence length and attention head variations on computational costs. When investigating the impact of sequence length, we set the attention head to 2 for EulerFormer and 4 for RecGRELA. When investigating the impact of attention heads, we set the sequence length to 100 to avoid GPU memory errors. Other hyperparameters are the same as the main experiments.
4.5.2. The Number of Sequence Length
@@
4.5.3. The Number of Attention Heads
@@
4.5.4. The Dropout and Droppath Rates
@@
Figure 8. Influence of the sequence length, head numbers, Dropout, and Droppath on various metrics.
4.6. Visualization Analysis
@@
Figure 9. Heatmap of output feature maps from the last layer.
5. Related Work
5.1. Sequential Recommendation
@@
Earlier fundamental models of sequential recommendation focus on using Convolutional Neural Networks (Tang and Wang, 2018; Yuan et al., 2020) and Recurrent Neural Networks (RNNs) (Hidasi, 2016; Li et al., 2017). Due to RNNs’ performance limits and poor parallel processing, recent sequential recommendation studies have increasingly turned to Transformer-based architectures (Wu et al., 2020; Li et al., 2021; Du et al., 2023; Fan et al., 2022; Zhang et al., 2023; Zhao et al., 2023; Li et al., 2023a; Shi et al., 2024) or MLPs (Long et al., 2024; Zhou et al., 2022; Li et al., 2023b, 2022; Gao et al., 2024; Jiang et al., 2024) for better efficiency and scalability. The most representative work is SASRec (Kang and McAuley, 2018), which is the first to use the self-attention mechanism for capturing user interest features from interaction sequences. BERT4Rec (Sun et al., 2019) adopts a deep encoder-based bidirectional model with self-attention to model user behavior sequences, achieving strong recommendation performance. However, the Transformer architecture suffers from quadratic computational complexity with respect to sequence length due to the dot-product attention operation, which limits its efficiency for sequence modeling tasks. Recently, many models with linear computational complexity have emerged as alternatives to the original Transformer architecture, such as linear attention (Liu et al., 2023; Tian et al., 2024), linear RNNs (Liu et al., 2024a; Yue et al., 2024), and state space models (Liu et al., 2024b; Qu et al., 2024; Fan et al., 2024). While LinRec (Liu et al., 2023) adopts linear attention to reduce the computational complexity of the dot-product attention, they still rely on learnable position encodings to capture positional information. EulerFormer (Tian et al., 2024) introduces a novel Euler transformation with a rotary form for unified modeling of both semantic and positional difference, yet uses the dot-product attention, which results in computational costs comparable to LinRec. To overcome these specific limitations, RecGRELA combines linear attention with rotary position encodings and introduces a gated mechanism to model both local and long-term user interests.
5.2. Linear Attention
@@
As opposed to dot-product attention, linear attention (Katharopoulos et al., 2020) replaces the Softmax with separate kernel functions, thereby reducing computational complexity to $\mathcal{O}(N)$ through a change in computation order. Though efficient, designing a linear attention module as effective as softmax attention is a nontrivial problem. Although linear attention has achieved some success in other domains, such as computer vision (Shen et al., 2021; Qin et al., 2022; Han et al., 2024a, b) and neural language processing (Wang et al., 2020; Qin et al., 2023; Yang et al., 2024). In sequential recommendation, despite existing linear attention-based methods (Fan et al., 2021; Liu et al., 2023) decreasing the computational complexity, they cannot effectively model position encoding within behavior sequence and fail to capture local preference in long-term interactions. This motivates us to design a more efficient and effective linear attention-based module to address these issues. In response, we present RecGRELA, which aims to realize the full potential of linear attention in sequential recommendation by overcoming these critical challenges in positional awareness and local preference modeling.
6. Conclusion and Future Work
@@
In this work, we propose RecGRELA, a new paradigm for long-term sequential recommendation. By integrating rotary position encoding into a linear attention mechanism, our Rotary-Enhanced Linear Attention module captures long-range dependencies with reduced computational complexity. To model local interaction in long-range dependencies, we design a Local Shortcut operation and provide theoretical insight from the perspective of the rank of the matrix. To address the difficulty of the model in perceiving and adapting to dynamic user preference shifts, we propose a SiLU-based gating mechanism that dynamically balances local and long-term representations. Extensive experiments on four benchmark datasets show that RecGRELA outperforms state-of-the-art methods based on RNNs, Transformers, and Mamba architectures, achieving superior performance with significantly lower memory overhead.
@@
Acknowledgements.
This work is funded by the National Natural Science Foundation of China (Grant No. 72074036, 62072060) and supported by the Natural Science Foundation of Chongqing, China, No. CSTB2024NSCQ-MSX0701, CSTB2024NSCQ-MSX0617.

### v2

@@ 3.4 Prediction Layer and Training
$\hat{\mathcal{Y}}=\operatorname{Softmax}(\boldsymbol{E}^{\top}\boldsymbol{H}_{L}).$ (24)
Hence, following previous works [19, 30], we expect to minimize the cross-entropy between the predicted recommendation results $\hat{\mathcal{Y}}$, and the ground truth items $\mathcal{Y}$:
[…]
3.5 Complexity Analysis
@@
4 Experiments
Table 1: Statistics of the experimental datasets.
@@ 4 Experiments
4.1 Experimental Settings
4.1.1 Datasets
We conduct experiments and verify the effectiveness of our approach on five datasets, including MovieLens-1M, 20M, and 32M  [^1] https://grouplens.org/datasets/movielens/ (dubbed as ML-1M, ML-20M, ML-32M), LFM-1B  [^2] https://www.cp.jku.at/datasets/LFM-1b/, and Netflix  [^3] https://www.kaggle.com/datasets/netflix-inc/netflix-prize-data. The dataset statistics are shown in Table 1, where Avg. UA and Avg. IA denote the average actions of users and items. Notably, for MovieLens datasets, we filter out users and items with fewer than 5 interactions. For LFM-1B datasets, we use the processed atomic files of LFM1B-Artists (merged) and filter out users with fewer than 50 and more than 500 interactions. Netflix is a huge-scale dataset, containing approximately 100 million interactions and five thousand Avg. IA. To simplify the training process, we filter out users with fewer than 50 interactions while retaining items with interactions ranging from 50 to 15,000 for this dataset.
4.1.2 Baseline Models
To verify the effectiveness and efficiency of RecGRELA, we make comparisons between RecGRELA and strong RNN-, Attention-, and Mamba-based baselines: LinRec [31]: This model reduces the computational costs via linear attention in transformer-based models. We choose SASRec [19] as the backbone of LinRec.LRURec [62]: This method is based on linear recurrent units [36] for sequential recommendation. Mamba4Rec [30]: This is the first Mamba-based method, which explores applying the Mamba architecture in sequential recommendation tasks. EchoMamba4Rec [56]: This method processes sequences both forward and backward to enhance the sequential representation with Mamba and frequency domain filter. EulerFormer [50]: This method proposes a transformer variant with complex vectors based on Euler’s formula and a contrastive learning method to improve contextual representations. RecBLR [29]: This model shows a behavior-dependent linear recurrent unit to enhance user behavior modeling with a hardware-aware parallel scan approach. SIGMA [32]: This model employs a flipped Mamba to build a bidirectional framework to improve short sequence representations and contextual modeling.
Table 2: Performance comparisons on ML-1M, ML-20M, ML-32M, and Netflix datasets. Best and second-best results per row are shown in bold and underlined format, respectively. "Improv." denotes the percentage improvement of the best-performing model over the second-best for each metric. The $\star$” denotes the statistical significance ($p$ < 0.05) of the results of MegaRec compared to the best baseline.
@@ 4.1.2 Baseline Models
Figure 3: RecGRELA vs. baselines in terms of the GPU memory and FLOPs of the training stage on the ML-1M dataset.
4.1.3 Evaluation Protocols
@@
4.1.4 Implementation Details
We utilize the RecBole library  [^4] https://github.com/RUCAIBox/RecBole [66, 59], a unified framework for implementing sequential recommendations. To ensure a fair comparison, we fix the embedding size to 64, the hidden size of the MLP to 256, and the training batch size to 2048 for all the models. We directly adopt the values reported in their original manuscripts for the remaining hyperparameters of the baseline models. The number of blocks $L$ is set to 4. The number of attention heads $h$ is set to 4. The kernel size for 1D causal convolution is 4. We set the Dropout and Droppath rate to 0.1. We use Adam [21] with a learning rate of 0.001 for optimizing our model. The maximum sequence length $N$ is set to 200 for all the datasets. The early stopping epochs are 20 for ML-1M and 10 for other datasets. All the experiments are conducted on a single NVIDIA RTX4090.
4.2 Overall Performance
Table 2 reports the recommendation performance of RecGRELA compared to seven baselines on five datasets. From the table, we find several observations:
@@ 4.3 Efficiency Analysis
Table 3: Sensitivity analysis on different numbers of layers. The sequence length is set to 200. Time indicates the training time per epoch, and Memory indicates the GPU memory.
@@
Figure 4: The effectiveness of each variant of RecGRELA on ML-1M.
4.4 Ablation Study
4.4.1 Core Components
@@
Figure 5: The performance of different position encodings and activation functions.
4.4.2 Position Encoding and Activation Function
@@
Figure 6: Impact of different stacked layers on computational costs. We set the sequence length to 100 to avoid GPU memory errors.
4.5 Sensitivity Analysis
@@ 4.5.1 The Number of Layers
Figure 7: Impact of sequence length and attention head variations on computational costs. When investigating the impact of sequence length, we set the attention head to 2 for EulerFormer and 4 for RecGRELA. When investigating the impact of attention heads, we set the sequence length to 100 to avoid GPU memory errors. Other hyperparameters are the same as the main experiments.
4.5.2 The Number of Sequence Length
@@
4.5.3 The Number of Attention Heads
@@
4.5.4 The Dropout and Droppath Rates
@@
Figure 8: Influence of the sequence length, head numbers, Dropout, and Droppath on various metrics.
Figure 9: Heatmap of output feature maps from the last layer.
4.6 Visualization Analysis
@@
5 Related Work
5.1 Sequential Recommendation
@@
Earlier fundamental models of sequential recommendation focus on using Convolutional Neural Networks [49, 61] and Recurrent Neural Networks (RNNs) [18, 25]. Due to RNNs’ performance limits and poor parallel processing, recent sequential recommendation studies have increasingly turned to Transformer-based architectures [57, 28, 5, 8, 64, 65, 24] or MLPs [33, 67, 27] for better efficiency and scalability. The most representative work is SASRec [19], which is the first to use the self-attention mechanism for capturing user interest features from interaction sequences. BERT4Rec [47] adopts a deep encoder-based bidirectional model with self-attention to model user behavior sequences, achieving strong recommendation performance. However, the Transformer architecture suffers from quadratic computational complexity with respect to sequence length due to the dot-product attention operation, which limits its efficiency for sequence modeling tasks. Recently, many models with linear computational complexity have emerged as alternatives to the original Transformer architecture, such as linear attention [31, 50], linear RNNs [29, 62], and state space models [30, 40, 6]. While LinRec [31] adopts linear attention to reduce the computational complexity of the dot-product attention, they still rely on learnable position encodings to capture positional information. EulerFormer [50] introduces a novel Euler transformation with a rotary form for unified modeling of both semantic and positional difference, yet uses the dot-product attention, which results in computational costs comparable to LinRec. To overcome these specific limitations, RecGRELA combines linear attention with rotary position encodings and introduces a gated mechanism to model both local and long-term user interests.
5.2 Linear Attention
@@
As opposed to dot-product attention, linear attention [20] replaces the Softmax with separate kernel functions, thereby reducing computational complexity to $\mathcal{O}(N)$ through a change in computation order. Though efficient, designing a linear attention module as effective as softmax attention is a nontrivial problem. Although linear attention has achieved some success in other domains, such as computer vision [44, 38, 14, 15] and neural language processing [55, 39, 60]. In sequential recommendation, despite existing linear attention-based methods [7, 31] decreasing the computational complexity, they cannot effectively model position encoding within behavior sequence and fail to capture local preference in long-term interactions. This motivates us to design a more efficient and effective linear attention-based module to address these issues. In response, we present RecGRELA, which aims to realize the full potential of linear attention in sequential recommendation by overcoming these critical challenges in positional awareness and local preference modeling.
6 Conclusion and Future Work
@@
In this work, we propose RecGRELA, a new paradigm for long-term sequential recommendation. By integrating rotary position encoding into a linear attention mechanism, our Rotary-Enhanced Linear Attention module captures long-range dependencies with reduced computational complexity. To model local interaction in long-range dependencies, we design a Local Shortcut operation and provide theoretical insight from the perspective of the rank of the matrix. To address the difficulty of the model in perceiving and adapting to dynamic user preference shifts, we propose a SiLU-based gating mechanism that dynamically balances local and long-term representations. Extensive experiments on five benchmark datasets show that RecGRELA outperforms state-of-the-art methods based on RNNs, Transformers, and Mamba architectures, achieving superior performance with significantly lower memory overhead.
@@
[…]